# FlexiTex: Enhancing Texture Generation with Visual Guidance

### v2

@@ Introduction
Figure 1: The framework of  FlexiTex. Given a mesh with corrected orientation and text/image prompts, the Visual Guidance Enhancement module (Sec. Visual Guidance Enhancement) extracts visual features to provide informative guidance during the denoising steps. The Direction-Aware Adaptation module (Sec. Direction-Aware Adaptation) then extracts direction features according to camera poses. Visual and direction features are integrated into the pre-trained UNet model with decoupled cross-attention, along with geometry features from depth maps. Through iterative texture warping and rasterization, our method generates high-fidelity textures of both rich details and multi-view consistency.
@@
FlexiTexis designed to generate high-quality texture maps given an untextured mesh conditioned on either text or image. The overview is shown in Fig. 1. In this section, we first provide the backgrounds of diffusion models, mesh rendering, and texture warping in Sec. Preliminary. Following this, we introduce the Visual Guidance Enhancement module in Sec. Visual Guidance Enhancement, which is designed to overcome issues related to over-smoothing. Then, we present the Direction-Aware Adaptation module in Sec. Direction-Aware Adaptation, which is specifically designed to incorporate direction information during texture generation, thereby addressing the Janus problem.

### v3

@@ Introduction
Figure 1: The framework of  FlexiTex. Given a mesh with corrected orientation and text/image prompts, the Visual Guidance Enhancement module extracts visual features to provide informative guidance during the denoising steps. The Direction-Aware Adaptation module then extracts direction features according to camera poses. Visual and direction features are integrated into the pre-trained UNet model with decoupled cross-attention, along with geometry features from depth maps. Through iterative texture warping and rasterization, our method generates high-fidelity textures of both rich details and multi-view consistency.
@@
FlexiTexis designed to generate high-quality texture maps given an untextured mesh conditioned on either text or image. The overview is shown in Fig. 1. In this section, we first provide the backgrounds of diffusion models, mesh rendering, and texture warping. Following this, we introduce the Visual Guidance Enhancement module, which is designed to overcome issues related to over-smoothing. Then, we present the Direction-Aware Adaptation module, which is specifically designed to incorporate direction information during texture generation, thereby addressing the Janus problem.